8x - 10y = -24
-8x + 6y = 40

Answer by tommyt3rd(5050)

D2: 8x - 10y = -24
D2: -8x + 6y = 40

4x-5y=-12
-4x+3y=20

-2y=8
y=-4


-8x+6(-4)=40

x=(40+24)/(-8)
x=-8

:)